### v1

@@ 3 Pooling-Based Methods
where the indicator variable $T_{i}$ is removed due to the perfect collinearity with $T_{i}\cdot\boldsymbol{\iota}_{L}(S_{i})$, since $\boldsymbol{\iota}_{L}(S_{i})$ contains indicator variables capturing a partitioning of $\mathcal{B}$, and $\otimes$ denotes the Kronecker product. The treatment effect of interest is now the collection of coefficients accompanying $T_{i}\cdot\boldsymbol{\iota}_{L}(S_{i})$, one for each segment. Specification (8) is roughly equivalent to employing specification (6) $L$ times, in each case only using the data corresponding to one of $L$ segments. (The only minor difference is whether observations are assigned to only one segment, or reused for different segment-specific treatment effect estimation implementation.) For instance, ou2010leave, LondonoVelezRodriguezSanchez_2020_AEJ and salti2022impact employ this boundary-segment specific pooling-based approach with $L=2$, analyzing each of the two linear segments as in Figure 1(b). Another empirical example employing this approach is .
Finally, some researchers have also included both the univariate distance $D_{i}$ and the bivariate score $\mathbf{X}_{i}$ in their local regression specifications—see, for example, ehrlich2018persistent and dehdari2022origins. In addition, it is possible to conceptualize more flexible specifications by including other interaction terms between $T_{i}$, $\mathbf{r}_{p}(\mathbf{X}_{i})$, and $\boldsymbol{\iota}_{L}(S_{i})$ analogous to specification (8). These approaches are uncommon in practice according to our literature review (Table 5), so we do not discuss them further.
@@ 3.1 Methodological Results
While there has been a proliferation of different empirical strategies relaying on the pooling of observations closest to the treatment assignment boundary, little is known about their properties and relative merits for the analysis of BD designs. Cattaneo-Titiunik-Yu_2025_BDD-Pooling recently investigated the main empirical approaches employing pooling-based methods, offering formal identification, estimation, and inference results. Their methodological results rely on techniques from geometric measure theory (federer2014geometric), which are used to characterize how the different treatment effect estimates emerging from specifications (1) through (6) behave in large samples, that is, when $h\to 0$ as $n\to 0$. These asymptotic approximations form the basis for the continuity-based approach in standard univariate RD designs (Hahn-Todd-vanderKlaauw_2001_ECMA), and are nowadays the most common way of characterizing the properties of causal treatment effect estimators in all RD settings. This section summarizes some of the findings in Cattaneo-Titiunik-Yu_2025_BDD-Pooling.
@@
The main challenge when studying pooling-based methods for BD designs is that the localization is done using a sequence of shrinking tubular neighborhoods covering the one-dimensional submanifold $\mathcal{B}$, as illustrated in Figure 2. An important implication is that the geometry of the treatment assignment boundary and the specific distance function used can substantially affect statistical properties. In this section, we omit most technicalities, and focus instead on the main ideas and conclusions to aid empirical researchers interpret and implement pooling-based methods.
@@
Because the treatment assignment boundary $\mathcal{B}$ could be highly irregular, particularly in Geographic RD designs, it is necessary to restrict its geometry to establish formal properties. A minimal assumption commonly used in mathematics is to require $\mathcal{B}$ to be a rectifiable curve, that is, to assume that it has finite length. More formally, $\mathcal{B}$ is a rectifiable curve if and only if it can be reparameterized by a Lipschitz continuous function. This restriction allows for formally defining and computing integrals of functions along the one-dimensional domain $\mathcal{B}$. More importantly, under additional technical conditions, regularity on the geometry of $\mathcal{B}$ allows for those integrals to be computed over shrinking tubular neighborhoods and for limits to be well-defined, which is at the core of the results in Cattaneo-Titiunik-Yu_2025_BDD-Pooling. Under regularity conditions, the authors establish the following result:
@@
$\lim_{h\downarrow 0}\int_{\mathcal{T}(h)}\frac{1}{h}g\Big(\frac{\mathcal{d}(\mathbf{x},\mathcal{B})}{h}\Big)m(\mathbf{x})\text{d}\mathbf{x}=\mathtt{c}_{\mathcal{B}}\cdot\int_{0}^{1}g(s)ds\cdot\int_{\mathcal{B}}\frac{m(\mathbf{x})}{\operatorname{J}_{2}\hskip-1.8063pt\mathcal{d}(\mathbf{x},\mathcal{B})}\text{d}\mathfrak{H}(\mathbf{x}),$ (9)
where $\mathcal{T}(h)=\big\{\mathbf{x}\in\mathcal{X}:\mathcal{d}(\mathbf{x},\mathcal{B})\leq h\big\}$, for each $h\geq 0$, is a tubular covering $\mathcal{B}$,
@@
and $\operatorname{J}_{2}\hskip-1.8063pt\mathcal{d}(\mathbf{x},\mathcal{B})$ denotes its Jacobian, and $\mathfrak{H}(\mathbf{x})$ denotes the $1$-dimensional Hausdorff measure. The constant $\mathtt{c}_{\mathcal{B}}$ can also be characterized, but it is not important for the upcoming discussion (it can be set to $\mathtt{c}_{\mathcal{B}}=2$ for the analysis of pooling-based methods). Note that $D_{i}=\big(\mathds{1}(\mathbf{X}_{i}\in\mathcal{A}_{1})-\mathds{1}(\mathbf{X}_{i}\in\mathcal{A}_{0})\big)\cdot\mathcal{d}(\mathbf{X}_{i},\mathcal{B})$.
The integral $\int_{\mathcal{B}}\frac{m(\mathbf{x})}{\operatorname{J}_{2}\hskip-1.8063pt\mathcal{d}(\mathbf{x},\mathcal{B})}\text{d}\mathfrak{H}(\mathbf{x})$ is a natural generalization of the standard line integral, where $\text{d}\mathfrak{H}$ serves as the rigorous “length element”, enabling a robust framework for integration over complex one-dimensional domains in the plane. In particular, if $\mathcal{B}$ is piecewise linear (as in Figures 1(b) and 2), the integral can be reduced to the sum of integrals over the linear segments using a natural smooth curve parametrization and standard Riemann integration (when $m(\mathbf{x})$ is Riemann integrable). For notational simplicity, we write
$\int_{\mathcal{B}}m(\mathbf{b})\text{d}\mathbf{x}=\int_{\mathcal{B}}\frac{m(\mathbf{x})}{\operatorname{J}_{2}\hskip-1.8063pt\mathcal{d}(\mathbf{x},\mathcal{B})}\text{d}\mathfrak{H}(\mathbf{x}),$
@@
The limiting integral in (9) can be used to give precise meaning to the probability limit of the treatment effect estimates from any of the pooling specifications discussed previously. However, we need to introduce standard potential outcomes notation (see, e.g., Hernan-Robins_2020_Book, for a review) to give a causal interpretation to the probability limit of those parameters. For $i=1,2,\dots,n$, let $Y_{i}(0)$ and $Y_{i}(1)$ denote the potential outcomes for unit $i$ under control and treatment assignment, respectively, and hence the observed outcome is
@@
To streamline the presentation, we consider only the simplest specification (1), leading to the local difference-in-means estimator $\widehat{\tau}_{\mathtt{DIM}}$, to illustrate the methodological findings in Cattaneo-Titiunik-Yu_2025_BDD-Pooling. The other more flexible local regression specifications will enjoy similar properties.
The probability limit of $\widehat{\tau}_{\mathtt{DIM}}$ can be naturally characterized as a ratio of integrals along the boundary. As $n\to\infty$ and leveraging the logic of the law of large numbers, $\frac{1}{nh}N_{1}(h)=\frac{1}{nh}\sum_{i=1}^{n}\mathds{1}(0\leq D_{i}\leq h)/(nh)$ will be close in probability to $\frac{1}{h}\mathbb{E}[\mathds{1}(0\leq D_{i}\leq h)]$, and
@@
The parameter $\tau$ is called the Boundary Average Treatment Effect (BATE) by Cattaneo-Titiunik-Yu_2025_BDD-Pooling. This parameter was heuristically introduced in the education literature by Wong-Steiner-Cook_2013_JEBS, who expressed it as
@@
using the notation $f(\mathbf{x}|\mathbf{X}_{i}\in\mathcal{B})=\frac{f(\mathbf{x})}{\int_{\mathcal{B}}f(\mathbf{x})\text{d}\mathbf{x}}$. Around the same time, Keele-Titiunik_2015_PA also discussed the BATE parameter in the context of Geographic RD designs.
The parameter $\tau$ captures a density-weighted average of average treatment effects at each point along the boundary, and thus aggregates the potentially heterogeneous treatment effects captured by $(\tau(\mathbf{x}):\mathbf{x}\in\mathcal{B})$. Cattaneo-Titiunik-Yu_2025_BDD-Location call $\tau(\mathbf{x})$ the Boundary Average Treatment Effect Curve. As discussed in Section 4, this functional causal parameter can be seen as a building block for several causal parameters of interest in BD designs.
Going beyond identification, Cattaneo-Titiunik-Yu_2025_BDD-Pooling also study estimation accuracy and uncertainty quantification for pooling-based methods. For point estimation, the authors shows that
@@
when $h\to 0$ as $n\to\infty$, and where $\approx_{\mathbb{P}}$ denotes up to higher order terms in probability, $\mathsf{B}_{n}$ denotes a leading bias term, and $\mathsf{V}$ denotes the asymptotic variance that is bounded and bounded away from zero. Characterizing the precise order of the bias is difficult without additional restrictions on the treatment assignment boundary $\mathcal{B}$: the authors show that $\mathsf{B}_{n}=O(h)$ in general. Furthermore, as in the case of standard univariate RD designs, and under additional regularity conditions, the order of the bias may be improved when $\mathbf{r}_{p}(D_{i})$ and $T_{i}\cdot\mathbf{r}_{p}(D_{i})$ are included in the local regression specification, as in (6), leading to $\mathsf{B}_{n}=O(h^{p+1})$. Another interesting feature is that the rate of decay of the variance component is the same as in one-dimensional nonparametric estimation: this improvement in convergence rate comes from the aggregation along the boundary of the bivariate nonparametric function $\tau(\mathbf{x})$. See Cattaneo-Titiunik-Yu_2025_BDD-Location and chen2025semiparametric for more discussion.
For bandwidth selection, under regularity conditions on $\mathcal{B}$ and $\mathcal{d}(\cdot)$, the bias $\mathsf{B}_{n}$ may admit a valid expansion as in the standard RD design literature. Therefore, optimal bandwidth selection procedures based on mean square error (MSE) or coverage error minimization may be used for implementation of pooling methods. See Calonico-Cattaneo-Farrell_2020_ECTJ for a review.
@@
where $\widehat{\mathbb{V}}[\widehat{\tau}_{\mathtt{DIM}}]$ denotes any of the usual variance estimators from (local, weighted) least squares regression methods. Under regularity conditions, and if $nh\to\infty$ and $nh\mathsf{B}_{n}^{2}\to 0$, $\mathbb{P}[\widehat{\operatorname{T}}\leq u]\to\Phi(u)$, where $\Phi(u)$ denotes the standard Gaussian cumulative distribution function. Importantly, as discussed by Calonico-Cattaneo-Titiunik_2014_ECMA in the context of standard univariate RD designs, the usual confidence intervals
@@
with $\mathfrak{c}_{\alpha}=\Phi^{-1}(1-\alpha/2)$, will not be valid whenever there is local misspecification. More precisely, the validity of the usual confidence intervals $\widehat{\operatorname{I}}(\alpha)$ require the small bias condition $nh\mathsf{B}_{n}^{2}\to 0$, which is not satisfied when the bandwidth is chosen to be MSE-optimal (or is otherwise “large”). Therefore, valid uncertainty quantification depends on the choice of bandwidth and regression specification, because the bias $\mathsf{B}_{n}$ would be of order $h^{p}$, under regularity conditions on $\mathcal{B}$, $\mathcal{d}(\cdot)$, and the underlying conditional expectation functions.
If the MSE-optimal bandwidth is used, which in the context of pooling-based methods would depend on the order of the bias $\mathsf{B}_{n}$, then the small bias condition will not be satisfied, and thus the resulting inference procedures will over-reject the null hypothesis, sometimes substantially. Following Calonico-Cattaneo-Titiunik_2014_ECMA, a simple and practical solution is to employ robust bias correction. The two-step procedure is simple and intuitive: first, the pooled treatment effect is estimated using a choice of polynomial approximation $p$ and its associated MSE-optimal bandwidth; second, confidence intervals and hypothesis tests are constructed by estimating the pooled treatment effect and variance estimator using the same bandwidth choice but with a polynomial of larger order $q$. In practice, the most common choice is $p=1$ for MSE-optimal treatment effect estimation, and $q=2$ for robust bias-corrected uncertainty quantification. This inference approach has several theoretical advantages (Calonico-Cattaneo-Farrell_2018_JASA; Calonico-Cattaneo-Farrell_2022_Bernoulli), and has been validated empirically (Hyytinen-Tukiainen-etal2018_QE; DeMagalhaes-etal_2025_PA). Standard RD software for estimation and inference in univariate score settings (rdrobust, rdhte) available at https://rdpackages.github.io/ can be used directly, provided the regularity conditions are satisfied.

### v2

@@ 3 Pooling-Based Methods
where the indicator variable $T_{i}$ is removed due to the perfect collinearity with $T_{i}\cdot\boldsymbol{\iota}_{L}(S_{i})$, since $\boldsymbol{\iota}_{L}(S_{i})$ contains indicator variables capturing a partitioning of $\mathcal{B}$, and $\otimes$ denotes the Kronecker product. The treatment effect of interest is now the collection of coefficients accompanying $T_{i}\cdot\boldsymbol{\iota}_{L}(S_{i})$, one for each segment. Specification (8) is roughly equivalent to fitting specification (6) $L$ times, in each case only using the data corresponding to one of $L$ segments—the only difference is whether observations are assigned to a single segment or reused). Ou (2010), Londoño-Vélez et al. (2020) and Salti et al. (2022) employ this boundary-segment specific pooling-based approach with $L=2$, analyzing each of the two linear segments as in Figure 1(b).
Some researchers have also included both the univariate distance $D_{i}$ and the bivariate score $\mathbf{X}_{i}$ in their local regression specifications—see, for example, Ehrlich and Seidel (2018) and Dehdari and Gehring (2022). It is also possible to conceptualize more flexible specifications by including other interaction terms between $T_{i}$, $\mathbf{r}_{p}(\mathbf{X}_{i})$, and $\boldsymbol{\iota}_{L}(S_{i})$ analogous to specification (8). Our literature review in Table 5 suggests that these approaches are uncommon in practice, so we do not discuss them further.
Although we have not systematically categorized the bivariate score in each of the empirical applications in Table 5, we suspect that some of the specification choices in practice may be determined by data availability and the granularity of the score. For example, Dell (2010) uses latitude and longitude of districts as the bivariate score, but her unit of observation is individuals, creating equal score values or mass points for all individuals in the same district (mass points). When limited data availability prevents the collection of continuously distributed bivariate scores, the statistical properties and methods we discuss in the upcoming sections do not apply directly, as they all assume that the score is a continuously distributed random variable. Keele and Titiunik (2015, 2016) discuss the importance of score granularity in geographic BD designs.
@@ 3.1 Methodological Results
While there has been a proliferation of different empirical strategies relying on the pooling of observations closest to the treatment assignment boundary, little is known about their properties and relative merits for the analysis of BD designs. Cattaneo et al. (2026) recently investigated pooling-based methods, offering formal identification, estimation, and inference results. Their methodological results rely on techniques from geometric measure theory (Federer, 2014), which are used to characterize how the different treatment effect estimates emerging from specifications (1) through (6) behave in large samples, that is, when $h\to 0$ as $n\to\infty$. These asymptotic approximations form the basis of the continuity-based approach in standard univariate RD designs (Hahn et al., 2001), and are nowadays the most common way of characterizing the properties of causal treatment effect estimators in all RD settings. This section summarizes some of the findings in Cattaneo et al. (2026).
@@
The main challenge of pooling-based BD methods is that the localization is done using a sequence of shrinking tubular neighborhoods covering the one-dimensional submanifold $\mathcal{B}$, as illustrated in Figure 2. An important implication is that the geometry of the treatment assignment boundary and the specific distance function used can substantially affect statistical properties. Our upcoming discussion omits most technicalities and focuses instead on conceptual ideas to aid empirical researchers interpret and implement pooling-based methods.
@@
Because the treatment assignment boundary $\mathcal{B}$ could be highly irregular, particularly in Geographic RD designs, it is necessary to restrict its geometry to establish formal properties. A minimal assumption commonly used in mathematics is to require $\mathcal{B}$ to be a rectifiable curve, that is, to assume that it has finite length (more formally, $\mathcal{B}$ is a rectifiable curve if and only if it can be reparameterized by a Lipschitz continuous function). This restriction allows for formally defining and computing integrals of functions along the one-dimensional domain $\mathcal{B}$. More importantly, under additional technical conditions, regularity on the geometry of $\mathcal{B}$ allows for those integrals to be computed over shrinking tubular neighborhoods and for limits to be well-defined, which is at the core of the results in Cattaneo et al. (2026). Under regularity conditions, the authors establish the following result:
@@
$\lim_{h\downarrow 0}\int_{\mathcal{T}(h)}\frac{1}{h}g\Big(\frac{\mathcal{d}(\mathbf{x},\mathcal{B})}{h}\Big)m(\mathbf{x})\text{d}\mathbf{x}=\mathtt{c}_{\mathcal{B}}\cdot\int_{0}^{1}g(s)ds\cdot\int_{\mathcal{B}}\frac{m(\mathbf{x})}{\operatorname{J}\hskip-1.8063pt\mathcal{d}(\mathbf{x},\mathcal{B})}\text{d}\mathfrak{H}(\mathbf{x}),$ (9)
where $\mathcal{T}(h)=\big\{\mathbf{x}\in\mathcal{X}:\mathcal{d}(\mathbf{x},\mathcal{B})\leq h\big\}$, for each $h\geq 0$, is a tubular neighborhood covering $\mathcal{B}$,
@@
and $\operatorname{J}\hskip-1.8063pt\mathcal{d}(\mathbf{x},\mathcal{B})$ denotes its Jacobian, and $\mathfrak{H}(\mathbf{x})$ denotes the $1$-dimensional Hausdorff measure. The constant $\mathtt{c}_{\mathcal{B}}$ can also be characterized, but it is not important for the upcoming discussion (it can be set to $\mathtt{c}_{\mathcal{B}}=2$ for the analysis of pooling-based methods). Note that $D_{i}=\big(\mathds{1}(\mathbf{X}_{i}\in\mathcal{A}_{1})-\mathds{1}(\mathbf{X}_{i}\in\mathcal{A}_{0})\big)\cdot\mathcal{d}(\mathbf{X}_{i},\mathcal{B})$.
The integral $\int_{\mathcal{B}}\frac{m(\mathbf{x})}{\operatorname{J}\hskip-1.8063pt\mathcal{d}(\mathbf{x},\mathcal{B})}\text{d}\mathfrak{H}(\mathbf{x})$ is a natural generalization of the standard line integral, where $\text{d}\mathfrak{H}$ serves as the rigorous “length element”, enabling a robust framework for integration over complex one-dimensional domains in the plane. In particular, if $\mathcal{B}$ is piecewise linear (as in Figures 1(b) and 2), the integral can be reduced to the sum of integrals over the linear segments using a natural smooth curve parametrization and standard Riemann integration (when $m(\mathbf{x})$ is Riemann integrable). For notational simplicity, we write
$\int_{\mathcal{B}}m(\mathbf{b})\text{d}\mathbf{x}=\int_{\mathcal{B}}\frac{m(\mathbf{x})}{\operatorname{J}\hskip-1.8063pt\mathcal{d}(\mathbf{x},\mathcal{B})}\text{d}\mathfrak{H}(\mathbf{x}),$
@@
The limiting integral in (9) can be used to give precise meaning to the probability limit of the treatment effect estimates from any of the pooling specifications discussed previously. However, we need to introduce standard potential outcomes notation (see, e.g., Hernán and Robins, 2020) to give a causal interpretation to the probability limit of those parameters. For $i=1,2,\dots,n$, let $Y_{i}(0)$ and $Y_{i}(1)$ denote the potential outcomes for unit $i$ under control and treatment assignment, respectively. The observed outcome is
@@
To streamline the presentation, we illustrate the methodological findings in Cattaneo et al. (2026) using the simplest specification (1), leading to the local difference-in-means estimator $\widehat{\tau}_{\mathtt{DIM}}$. The other more flexible local regression specifications enjoy similar properties.
The probability limit of $\widehat{\tau}_{\mathtt{DIM}}$ can be naturally characterized as a ratio of integrals along the boundary. As $n\to\infty$ and leveraging the logic of the law of large numbers, $\frac{1}{nh}N_{1}(h)=\frac{1}{nh}\sum_{i=1}^{n}\mathds{1}(0\leq D_{i}\leq h)$ will be close in probability to $\frac{1}{h}\mathbb{E}[\mathds{1}(0\leq D_{i}\leq h)]$, and
@@
The parameter $\tau$ is called the Boundary Average Treatment Effect (BATE) by Cattaneo et al. (2026). This parameter was heuristically introduced in the education literature by Wong et al. (2013), who expressed it as
@@
using the notation $f(\mathbf{x}|\mathbf{X}_{i}\in\mathcal{B})=\frac{f(\mathbf{x})}{\int_{\mathcal{B}}f(\mathbf{x})\text{d}\mathbf{x}}$. Around the same time, Keele and Titiunik (2015) also discussed the BATE parameter in the context of Geographic RD designs.
The parameter $\tau$ captures a density-weighted average of average treatment effects at each point along the boundary, and thus aggregates the potentially heterogeneous treatment effects captured by $(\tau(\mathbf{x}):\mathbf{x}\in\mathcal{B})$. Cattaneo et al. (2025b) call $\tau(\mathbf{x})$ the Boundary Average Treatment Effect Curve (BATEC). As we discuss in Section 4, this functional causal parameter can be seen as a building block for several other causal parameters of interest in BD designs.
Going beyond identification, Cattaneo et al. (2026) study estimation accuracy and uncertainty quantification for pooling-based methods. For point estimation, the authors shows that
@@
when $h\to 0$ as $n\to\infty$, and where $\approx_{\mathbb{P}}$ denotes approximation in probability up to higher order terms, $\mathsf{B}_{n}$ denotes a leading bias term, and $\mathsf{V}$ denotes the asymptotic variance that is bounded and bounded away from zero. Characterizing the precise order of the bias is difficult without additional restrictions on the treatment assignment boundary $\mathcal{B}$: the authors show that $\mathsf{B}_{n}=O(h)$ in general. Furthermore, as in the case of standard univariate RD designs, and under additional regularity conditions, the order of the bias may be improved when $\mathbf{r}_{p}(D_{i})$ and $T_{i}\cdot\mathbf{r}_{p}(D_{i})$ are included in the local regression specification, as in (6), leading to $\mathsf{B}_{n}=O(h^{p+1})$. Another interesting feature is that the rate of decay of the variance component is the same as in one-dimensional nonparametric estimation: this improvement in convergence rate comes from the aggregation along the boundary of the bivariate nonparametric function $\tau(\mathbf{x})$. See Cattaneo et al. (2025b) and Chen and Gao (2025) for more discussion.
For bandwidth selection, under regularity conditions on $\mathcal{B}$ and $\mathcal{d}(\cdot)$, the bias $\mathsf{B}_{n}$ may admit a valid expansion as in the standard RD design literature. Therefore, optimal bandwidth selection procedures based on mean square error (MSE) or coverage error minimization may be used for implementation of pooling methods. See Calonico et al. (2020) for a review.
@@
where $\widehat{\mathbb{V}}[\widehat{\tau}_{\mathtt{DIM}}]$ denotes any of the usual variance estimators from (local, weighted) least squares regression methods. Under regularity conditions, and if $nh\to\infty$ and $nh\mathsf{B}_{n}^{2}\to 0$, $\mathbb{P}[\widehat{\operatorname{T}}\leq u]\to\Phi(u)$, where $\Phi(u)$ denotes the standard Gaussian cumulative distribution function. Importantly, as discussed by Calonico et al. (2014) in the context of standard univariate RD designs, the usual confidence intervals
@@
with $\mathfrak{c}_{\alpha}=\Phi^{-1}(1-\alpha/2)$, will not be valid whenever there is local misspecification. More precisely, the validity of the usual confidence intervals $\widehat{\operatorname{I}}(\alpha)$ require the small bias condition $nh\mathsf{B}_{n}^{2}\to 0$, which is not satisfied when the MSE-optimal bandwidth (or larger) is used. Therefore, the choice of bandwidth and regression specification determines whether uncertainty quantification is valid.
If the MSE-optimal bandwidth is used, which in the context of pooling-based methods would depend on the order of the bias $\mathsf{B}_{n}$, then the small bias condition will not be satisfied, and thus the resulting inference procedures will over-reject the null hypothesis, sometimes substantially. Following Calonico et al. (2014), a simple and practical solution is to employ robust bias correction. The two-step procedure is simple and intuitive: first, the pooled treatment effect is estimated using a choice of polynomial approximation $p$ and its associated MSE-optimal bandwidth; second, confidence intervals and hypothesis tests are constructed by estimating the pooled treatment effect and variance estimator using the same bandwidth choice but with a polynomial of larger order $q$. In practice, the most common choice is $p=1$ for MSE-optimal treatment effect estimation, and $q=2$ for robust bias-corrected uncertainty quantification. This inference approach has several theoretical advantages (Calonico et al., 2018, 2022), and has been validated empirically (Hyytinen et al., 2018; De Magalhães et al., 2025). Standard RD software for estimation and inference in univariate score settings (rdrobust, rdhte) available at https://rdpackages.github.io/ can be used directly, provided the regularity conditions are satisfied.